# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

### v4

@@ 3 Proofs
  • The symbol $\mathbb{E}$ represents the expectation operator associated with the probability space $(\Omega,\mathcal{F},\mathbb{P})$.
  • For any two Banach spaces $E_{1}$ and $E_{2}$, $\mathcal{L}(E_{1},E_{2})$ denotes the space of all bounded linear operators from $E_{1}$ to $E_{2}$, and $\mathcal{L}(E_{1},E_{1})$ is abbreviated to $\mathcal{L}(E_{1})$.
  • The operator $I$ denotes the identity map.
@@
  • For $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, $\dot{H}_{h}^{\alpha,q}$ is defined as the Banach space $X_{h}$ equipped with the norm
    $\lVert{v_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h}.$
  • For $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, equipped with the norm
    $\lVert{v}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}v}\rVert_{L^{q}},\quad\forall v\in\dot{H}^{\alpha,q}.$
  • Specifically, $\dot{H}^{-1,q}$ denotes the dual space of $\dot{H}^{1,q/(q-1)}$ for each $q\in(1,\infty)$.
  • The symbol $c$ is used to denote a generic positive constant, which is independent of $h$ and $\tau$, and its specific value may change from line to line.
The subsequent structure of this section is outlined as follows. In Subsection 3.1, we introduce several preliminary lemmas that form the foundation for our subsequent analysis. Subsequently, Subsection 3.2 is dedicated to establishing a stability estimate for a discrete stochastic convolution, which is of independent interest and also vital for the proof of Theorem 2.1. Lastly, Subsection 3.3 presents a detailed proof of Theorem 2.1.
@@ 3.1 Preliminary lemmas
For any $\alpha\in[0,2]$ and $q\in(1,\infty)$, $\lVert{P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\,\dot{H}_{h}^{\alpha,q})}$ is uniformly bounded with respect to (w.r.t.) the spatial mesh size $h$.
@@
Let $q\in(1,\infty)$ and $0\leqslant\alpha\leqslant\beta<\infty$. Then the following holds:
@@
Let $q\in[2,\infty)$. Then the following holds:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),$ $\quad\forall v\in L^{q},$ (3.5)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}),$ $\forall v\in\dot{H}^{1,q},$ (3.6)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ $\forall u,v\in L^{q}.$ (3.7)
@@
where $g\in L^{p}(0,T;L^{q})$. Then
@@
Remark 3.1.
Lemmas 3.1 and 3.2 are well known; see, e.g., [34]. For the proof of Lemma 3.3, we refer the reader to, e.g., [35, pp. 792]. Lemma 3.4 is contained in [22, Lemma 4.9]. For the proof of Lemma 3.5, we refer the reader to [16, 20].
@@ 3.2 A stability estimate for a discrete stochastic convolution
$\bigg{(}\mathbb{E}\Big{[}\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\bigg{)}^{1/p}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$ (3.8)
@@
$Z_{j}=\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}P_{h}g(t)\,\mathrm{d}W_{H}(t),\quad 1\leqslant j\leqslant J,$ (3.9)
where $g$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, with parameters $p\in(2,\infty)$ and $q\in[2,\infty)$. This stability estimate is crucial for the derivation of pathwise uniform convergence rates for the numerical approximations of a broad class of nonlinear stochastic parabolic equations.
Within the framework of Hilbert spaces, we note that Gyöngy and Millet [9, Theorem 4.4] have established the following stability estimate:
@@
Moreover, under the Banach space setting, a pivotal advancement was made by van Neerven and Veraar [36, Proposition 5.4], who demonstrated the following sharp estimate for a discrete stochastic convolution without spatial discretization:
@@
valid for all $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$ with $p,q\in[2,\infty)$. The derivation of this estimate relies on the contraction property of the operator $(I-\tau\Delta)^{-1}$ on $L^{q}$ for all $q\in[2,\infty)$. However, it is important to note that $(I-\tau\Delta_{h})^{-1}$ fails to possess the contraction property on $\dot{H}_{h}^{0,q}$ for any $q\in(2,\infty)$. Hence, Proposition 5.4 from [35] can not be directly applied to bound the discrete stochastic convolution defined by Eq. 3.9.
To address the aforementioned challenge, we employ the approximation property of $(I-\tau\Delta_{h})^{-1}$ to the identity operator $I$ as presented in Eq. 3.3, in conjunction with the discrete stochastic maximal $L^{p}$-regularity estimate for
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p},\quad p\in(2,\infty),\,q\in[2,\infty).$
We formally state the discrete stochastic maximal $L^{p}$-regularity estimate in the following proposition.
@@
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Assume that $(Z_{j})_{j=1}^{J}$ is defined by Eq. 3.9, where $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Then the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\Big{(}\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{)}^{1/p}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}.$
@@
Note that [23, Theorem 3.1] has demonstrated that the $H^{\infty}$-calculus of $-\Delta_{h}$ is uniformly bounded with respect to the spatial mesh size $h$. Subsequently, by closely following the arguments outlined in the proof of [21, Theorem 3.2], we establish the desired discrete stochastic maximal $L^{p}$-regularity estimate. ∎
@@
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Suppose that $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, and $(Z_{j})_{j=1}^{J}$ are defined by Eq. 3.9. Then the stability estimate Eq. 3.8 holds.
@@
$\Psi(t):=(I-\tau\Delta_{h})^{-1}P_{h}g(t),\quad\mathcal{G}(t):=p\sum_{j=0}^{J-1}\mathbbm{1}_{[t_{j},t_{j+1})}(t)\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\left\langle\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\Psi(t)\right\rangle,$
@@
$\hskip-195.00029ptI_{2}:=\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\Big{\lvert}{\int_{0}^{t}\mathcal{G}(t)\,\mathrm{d}W_{H}(t)}\Big{\rvert}\bigg{]},$
$I_{3}:=p(p-q)\sum_{j=0}^{J-1}\mathbb{E}\bigg{[}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\Big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\Big{\rangle}^{2}\,\mathrm{d}\theta\bigg{]},$
$\hskip-60.00009ptI_{4}:=p(q-1)\sum_{j=0}^{J-1}\mathbb{E}\bigg{[}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta\bigg{]}.$
@@
$\quad{}+p(p-q)\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta$
@@
$\max_{0\leqslant j\leqslant J}\left\|Z_{j}\right\|_{L^{q}}^{p}\leqslant p\sum_{j=0}^{J-1}\left\|Z_{j}\right\|_{L^{q}}^{p-q}\left|\left\langle\left|Z_{j}\right|^{q-2}Z_{j},\delta_{j}\right\rangle\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\mathcal{G}(s)\,\mathrm{d}W_{H}(s)\right|$
$\qquad\qquad{}+p(p-q)\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-2q}\left\langle\left|Z_{j}+\theta M_{j}\right|^{q-2}(Z_{j}+\theta M_{j}),M_{j}\right\rangle^{2}\,\mathrm{d}\theta$
$\qquad\qquad{}+p(q-1)\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-q}\left\langle\left|Z_{j}+\theta M_{j}\right|^{q-2},M_{j}^{2}\right\rangle\,\mathrm{d}\theta,$
which directly implies the desired inequality (3.10) upon taking expectations on both sides.
Step 2. Let us prove
$I_{1}\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{]}.$ (3.11)
From the definition of $I_{1}$, it follows that
$I_{1}$ $\leqslant p\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}^{-1,q}}\Big{]}$
$\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}_{h}^{-1,q}}\Big{]}.$
@@
the last inequality justified by (3.3) with $\alpha=0$ and $\beta=2$. Furthermore,
@@
$\leqslant c\|Z_{j}\|_{L^{q}}^{q-2}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}.$
Consequently, we arrive at
$I_{1}\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\|Z_{j}\|_{L^{q}}^{p-2}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}^{2}\Big{]},$
which leads to Eq. 3.11 by Young’s inequality.
Step 3. Let us prove
$I_{2}\leqslant\varepsilon\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}+c\varepsilon^{1-p}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},\quad\forall\varepsilon\in(0,1).$ (3.12)
By the Burkholder-Davis-Gundy inequality and Hölder’s inequality, we obtain
@@
For any $\varepsilon\in(0,1)$, we can further bound $I_{2}$ as follows:
$I_{2}$ $\leqslant c\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p-1}\lVert{\Psi}\rVert_{L^{2}(0,T;\gamma(H,L^{q}))}\Big{]}$
$\leqslant c\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p-1}\lVert{\Psi}\rVert_{L^{p}(0,T;\gamma(H,L^{q}))}\Big{]}$
$\leqslant\varepsilon\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}+c\varepsilon^{1-p}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
where the second inequality uses Hölder’s inequality, and the third inequality applies Young’s inequality. This proves the desired inequality Eq. 3.12.
Step 4. Let us prove that
$I_{3}+I_{4}\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$ (3.13)
To this end, we apply Hölder’s inequality and Young’s inequality accordingly:
$I_{3}+I_{4}$ $\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2}\lVert{M_{j}}\rVert_{L^{q}}^{2}\,\mathrm{d}\theta$
$\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\int_{0}^{1}\tau\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p}+\tau^{1-p/2}\lVert{M_{j}}\rVert_{L^{q}}^{p}\,\mathrm{d}\theta$
$\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+c\tau^{1-p/2}\sum_{j=0}^{J-1}\mathbb{E}\lVert{M_{j}}\rVert_{L^{q}}^{p}.$
On the other hand,
$\sum_{j=0}^{J-1}\mathbb{E}\lVert{M_{j}}\rVert_{L^{q}}^{p}$ $\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\sum_{j=0}^{J-1}\mathbb{E}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}+c\sum_{j=0}^{J-1}\mathbb{E}\big{\lVert}{\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t)}\big{\rVert}_{L^{q}}^{p}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\sum_{j=0}^{J-1}\mathbb{E}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}+c\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\int_{t_{j}}^{t_{j+1}}\lVert{\Psi(t)}\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\Big{]}^{p/2}$

### v5

@@ 4 Proofs
  • For $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, $\dot{H}_{h}^{\alpha,q}$ represents the Banach space $X_{h}$ equipped with the norm
    $\lVert{u_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}u_{h}}\rVert_{L^{q}},\quad\forall u_{h}\in X_{h}.$
  • The symbol $c$ denotes a generic positive constant, which is independent of the spatial mesh size $h$ and the time step $\tau$. Its specific value may depend on the regularity parameters associated with the triangulation $\mathcal{K}_{h}$, and may change from line to line.
The subsequent structure of this section is outlined as follows. In Subsection 4.1, we introduce several preliminary lemmas that form the foundation for our subsequent analysis. Subsequently, Subsection 4.2 is dedicated to establishing a stability estimate for a discrete stochastic convolution, which is of independent interest and also crucial for the proof of Theorem 3.1. Lastly, Subsection 4.3 presents a detailed proof of Theorem 3.1.
@@ 4.1 Preliminary lemmas
According to the theoretical results in [9], the projection operator $P_{h}$ has the following well-known stability property.
@@
For any $\alpha\in[0,2]$ and $q\in(1,\infty)$, $\lVert{P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\,\dot{H}_{h}^{\alpha,q})}$ is uniformly bounded with respect to the spatial mesh size $h$.
Referring to the established estimates for the resolvents of $\Delta_{h}$ as established in [2], a standard computation (for which the relevant techniques can be found in [44, Section 7.7 in Chapter 2]) yields the following well-recognized inequalities.
@@
Let $q\in(1,\infty)$ and $0\leqslant\alpha\leqslant\beta<\infty$. Then the following inequalities hold:
@@
We have the following fundamental estimate for the stochastic integrals; see, e.g., [42, pp. 792].
@@
Suppose $q\in[2,\infty)$. The following inequalities hold:
$\lVert{F(u)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{u}\rVert_{L^{q}})$ for all $\text{for all }u\in L^{q},$ (4.5)
$\lVert{F(u)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{u}\rVert_{\dot{H}^{1,q}})$ for all $\text{for all }u\in\dot{H}^{1,q},$ (4.6)
$\lVert{F(u_{1})-F(u_{2})}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u_{1}-u_{2}}\rVert_{L^{q}}$ for all $\text{for all }u_{1},u_{2}\in L^{q}.$ (4.7)
Proof.
These properties are derived from straightforward calculations based on the definition of $F$. For a detailed derivation, see Lemma 4.5 in [28]. ∎
We have the following discrete deterministic maximal $L^{p}$-regularity estimate; see [20, 26].
@@
where $g\in L^{p}(0,T;L^{q})$. Then the following stability estimate holds:
@@
Finally, we present some standard properties of the space $\dot{H}_{h}^{\alpha,q}$, where $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$. For any $q\in(1,\infty)$, there exist two positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that for all $u_{h}\in X_{h}$, the following inequality holds:
$c_{0}\lVert{u_{h}}\rVert_{\dot{H}^{1,q}}\leqslant\lVert{u_{h}}\rVert_{\dot{H}_{h}^{1,q}}\leqslant c_{1}\lVert{u_{h}}\rVert_{\dot{H}^{1,q}}.$
For any $\alpha\in(0,2)$ and $q\in(1,\infty)$, there exist two positive constants $c_{2}$ and $c_{3}$, independent of $h$, such that for all $u_{h}\in X_{h}$, the following inequality holds:
$c_{2}\lVert{u_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}\leqslant\lVert{u_{h}}\rVert_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{3}\lVert{u_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}},$
where $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ is the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [44, Section 5 in Chapter 1]). Furthermore, for any $\beta\in[3/2,\infty)$ and $q\in[2,\infty)$, using the well-known estimate (see, e.g., [6, Section 8.5]),
$\lVert{\Delta^{-1}-\Delta_{h}^{-1}P_{h}}\rVert_{\mathcal{L}(L^{\beta})}\leqslant ch^{2},$
combined with the continuous embedding of $\dot{H}^{2,\beta}$ into $L^{q}$ and the inverse estimate, it follows that the continuous embedding of $\dot{H}_{h}^{2,\beta}$ into $L^{q}$ is independent of $h$. These results will be used throughout the subsequent analysis without further explicit reference.
@@ 4.2 A stability estimate for a discrete stochastic convolution
$\bigg{(}\mathbb{E}\Big{[}\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\bigg{)}^{1/p}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}$ (4.8)
@@
$Z_{j}=\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}g_{h}(t)\,\mathrm{d}W_{H}(t),\quad 1\leqslant j\leqslant J,$ (4.9)
where $g_{h}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, with parameters $p\in(2,\infty)$ and $q\in[2,\infty)$. This stability estimate is crucial for the derivation of pathwise uniform convergence rates for the numerical approximations of a broad class of nonlinear stochastic parabolic equations.
Within the framework of Hilbert spaces, we note that Gyöngy and Millet [13, Theorem 2.6] have established the following stability estimate:
@@
Moreover, under the Banach space setting, a significant advancement was made by van Neerven and Veraar [40, Proposition 5.4], who demonstrated the following sharp estimate for a discrete stochastic convolution without spatial discretization:
@@
holds for all $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$ with $p,q\in[2,\infty)$. The derivation of this estimate relies on the contraction property of the operator $(I-\tau\Delta)^{-1}$ on $L^{q}$ for all $q\in[2,\infty)$. However, it is important to note that $(I-\tau\Delta_{h})^{-1}$ fails to possess the contraction property on $\dot{H}_{h}^{0,q}$ for any $q\in(2,\infty)$. Hence, Proposition 5.4 from [42] can not be directly applied to bound the discrete stochastic convolution defined by Eq. 4.9.
To address the aforementioned challenge, we employ the discrete stochastic maximal $L^{p}$-regularity estimate for
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p},\quad p\in(2,\infty),\,q\in[2,\infty),$
coupled with the approximation property of $(I-\tau\Delta_{h})^{-1}$ to the identity operator $I$ as presented in Eq. 4.3. We formally state the discrete stochastic maximal $L^{p}$-regularity estimate in the following proposition.
@@
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Assume that $(Z_{j})_{j=1}^{J}$ is defined by Eq. 4.9, where $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$. Then the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\Bigg{(}\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}\Bigg{)}^{1/p}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}.$
@@
Note that [29, Theorem 3.1] has demonstrated that the $H^{\infty}$-calculus of $-\Delta_{h}$ is uniformly bounded with respect to the spatial mesh size $h$. Subsequently, applying [27, Theorem 3.2], we establish the desired discrete stochastic maximal $L^{p}$-regularity estimate. ∎
@@
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Suppose that $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, and let $(Z_{j})_{j=1}^{J}$ be defined by Eq. 4.9. Then the stability estimate Eq. 4.8 holds.
@@
$\Psi(t):=(I-\tau\Delta_{h})^{-1}g_{h}(t),\quad\mathcal{G}(t):=p\sum_{j=0}^{J-1}\mathbbm{1}_{[t_{j},t_{j+1})}(t)\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\left\langle\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\Psi(t)\right\rangle,$
@@
$\hskip-195.00029ptI_{2}:=\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\Big{\lvert}{\int_{0}^{t}\mathcal{G}(s)\,\mathrm{d}W_{H}(s)}\Big{\rvert}\bigg{]},$
$I_{3}:=p|p-q|\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta\bigg{]},$
$\hskip-60.00009ptI_{4}:=p(q-1)\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta\bigg{]}.$
@@
$\quad{}+p|p-q|\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta$
@@
$\max_{0\leqslant j\leqslant J}\left\|Z_{j}\right\|_{L^{q}}^{p}\leqslant p\sum_{j=0}^{J-1}\left\|Z_{j}\right\|_{L^{q}}^{p-q}\left|\big{\langle}{\left|Z_{j}\right|^{q-2}Z_{j},\,\delta_{j}}\big{\rangle}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\mathcal{G}(s)\,\mathrm{d}W_{H}(s)\right|$
$\qquad\qquad{}+p|p-q|\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-2q}\big{\langle}{\left|Z_{j}+\theta M_{j}\right|^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta$
$\qquad\qquad{}+p(q-1)\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-q}\big{\langle}{\left|Z_{j}+\theta M_{j}\right|^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta.$
By taking expectations on both sides of the inequality, we arrive at the desired inequality (4.10).
Step 2. From the definition of $I_{1}$, it follows that
$I_{1}$ $\leqslant p\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}^{-1,q}}\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}_{h}^{-1,q}}\bigg{]}.$
@@
where the last inequality is justified by (4.3) with $\alpha=0$ and $\beta=2$. Furthermore,
@@
$\leqslant c\|Z_{j}\|_{\dot{H}_{h}^{1,q}}^{q-1}.$
Consequently, putting the above estimates together, we obtain
$I_{1}\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}.$ (4.11)
Step 3. By the Burkholder-Davis-Gundy inequality and Hölder’s inequality, we obtain
@@
It follows that
$I_{2}\leqslant c\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p-1}\lVert{\Psi}\rVert_{L^{2}(0,T;\gamma(H,L^{q}))}\Big{]}.$
By Young’s inequality, we then obtain
$I_{2}\leqslant\frac{1}{2}\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))}^{p}.$ (4.12)
Step 4. We employ Hölder’s inequality and Young’s inequality as follows:
$I_{3}+I_{4}$ $\leqslant c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2}\lVert{M_{j}}\rVert_{L^{q}}^{2}\,\mathrm{d}\theta\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\int_{0}^{1}\tau\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p}+\tau^{1-p/2}\lVert{M_{j}}\rVert_{L^{q}}^{p}\,\mathrm{d}\theta\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\bigg{]}+c\tau^{1-p/2}\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{M_{j}}\rVert_{L^{q}}^{p}\bigg{]}.$
Moreover,